wiki:Research/LhARA/DesignAndIntegration/Meetings/2023/2023-08-22

Design and Integration meeting: 22 August 2023 15:00 GMT

Zoom: https://ukri.zoom.us/j/99571436156

Agenda

  • Actions from last meeting.
  • Simulations.
  • Engineering.
  • AOB

Actions

23-01-24-01 All to prepare document for the milestone to update the baseline In progress

Notes

Present: Neil Bliss, Clive Hill, Ajit Kurup, Ken Long, Jaroslaw Pasternak, Rehanah Razak, Will Shields, Colin Whyte
Apologies: Mark Johnson, Hywel Owen

  • Simulations
    • Update from JP:
      • Updated diagram of the ring.
        • Need to add details of the yoke.
      • Contour plot of Total B vs R and opening angle. Plot of B vs opening angle.
      • Need to add the injection line
    • Slides from WS.
      • Update on optics optimisation.
      • Lilli Platt has now finished.
      • GPT solution does not converge for spot sizes sigma=2cm or less.
      • Discussion with Tony Price, suggested smaller spot sizes not needed.
      • KL: We should provide smaller spots sizes but not necessary to deliver in the beam line, can use final focussing elements in the end station.
      • CW: Also having discussions with Tony regarding what measurements should be done for the end station.
      • Discussion on what should go in the 12 month report.
        • Should reference a technical note.
        • NB: Better to add the detail directly into the report for visibility.
  • Engineering
    • Update on the schematic based on WS's comment.
      • Ross Gray is keen on the vertical arrangement of the laser.
      • Investigating having the laser beam at 25 degrees following suggestion from Ross.
      • Discussion on effect of the beam energy due to different incident angle and flexibility for different target production mechanisms.
      • NB would like comments on the schematic.
      • JP: FETS uses wire scanners for profile measurements.
      • Discussion on profile monitor technology.

  • Reference design for the bunching cavities.
    • JP: Calculated the details of the cavity. Need to dig these out.
  • Schematic from CH showing the 25 degree target angle. Incoming laser angle is now 40 degrees.
  • Meeting tomorrow on power consumption.

Last modified 6 months ago Last modified on Aug 22, 2023 9:36:45 PM

Attachments (4)

Download all attachments as: .zip